OpenStudy (sleepyjess):

Find all solutions in the interval \([0, 2\pi)\) \(\sf 7 tan^3x - 21 tan x = 0\)

OpenStudy (sleepyjess):

@sammixboo

sammixboo (sammixboo):

Yeah I am not that good with pre-calc and I can't help with this. Some other stuff - I can. So sorry!

OpenStudy (sleepyjess):

:(

sammixboo (sammixboo):

Maybe @ganeshie8 can help?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{7 tan^3x - 21 tan x = 0\\~\\ 7 tanx (tan^2x- 3 ) = 0\\~\\ 7 tanx = 0~~\And~~(tan^2x- 3 ) =0\\~\\ x = arctan(0)~~\And~~tan^2x =3\\~\\ x = 0^{\circ}~~\And~~tanx =\pm \sqrt{3}\\~\\ x = 0^{\circ}~~\And~~x =\pm \dfrac{\pi}{3}\\~\\ }\end{align}\)

OpenStudy (anonymous):

does that make sense? @sleepyjess

OpenStudy (sleepyjess):

No, not really

OpenStudy (anonymous):

What part don't you understand?

OpenStudy (anonymous):

\(\large \begin{align} \color{black}{7 \tan^3x - 21 \tan x = 0\\~\\ \small\text{Factor 7 tanx out:}\\ 7 \tan x (\tan^2x- 3 ) = 0\\~\\ \small\text{We can see that either }7\tan x = 0 ~\text{ or }~ \tan^2x-3 = 0\text{, so}\\ 7 \tan x = 0~~\text{or} ~~(tan^2x- 3 ) =0\\~\\ x = \arctan(0)~~\text{or}~~\tan^2x =3\\~\\ x = 0^{\circ}~~\text{or}~~\tan x =\pm \sqrt{3}\\~\\ x = 0^{\circ}~~\text{or}~~x =\pm \dfrac{\pi}{3}\\~\\ }\end{align}\) Does that help at all?

OpenStudy (sleepyjess):

Where does \(\tan^2x\) come from?

OpenStudy (anonymous):

We started with \(7\tan^3x-21\tan x\) . We factored 7 tan x out. Like \(7u^3-21u\Longrightarrow7u ~(u^2-3)\)

OpenStudy (sleepyjess):

Oh ok

OpenStudy (anonymous):

ok, so everything makes sense now?

OpenStudy (sleepyjess):

I think so, but that isn't one of the options.

OpenStudy (anonymous):

interval is \([0, 2\pi)\) So answer is actually \(x=0~~\text{or}~~x = \dfrac{\pi}{3}\)

OpenStudy (anonymous):

hmm yeah something is missing

OpenStudy (sleepyjess):

Here are the options, http://prntscr.com/5h4skx

OpenStudy (anonymous):

@mathmath333 made a mistake. \(\tan x = \pm\sqrt3\cancel\longrightarrow x = \pm\dfrac{\pi}{3}\) if \(\tan x = \pm\sqrt3\), then either \(\tan x = \sqrt3\) or \(\tan x = -\sqrt3\) So we have \(\tan x = \sqrt3\) at \(x = \dfrac{\pi}{3}\) and \(x = \dfrac{4\pi}{3}\) and we have \(\tan x = -\sqrt3\) at \(x = \dfrac{2\pi}{3}\) and \(x = \dfrac{5\pi}{3}\) So that would be option C

OpenStudy (anonymous):

Oh, and this: \(\tan x=0~~\longrightarrow~~x=0\) or \(x = \pi\)

OpenStudy (anonymous):

you have the following solutions: \(x=0, \dfrac{\pi}{3}, \dfrac{2\pi}{3}, \dfrac{3\pi}{3}=\pi, \dfrac{4\pi}{3}, \dfrac{5\pi}{3}\)

OpenStudy (sleepyjess):

Thanks for explaining that to me. Can you help on a couple more?

OpenStudy (anonymous):

I have to go, sorry. Start a new question.

OpenStudy (sleepyjess):

Ok, thanks